# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v1

@@ Abstract
Sparse attention can effectively mitigate the significant memory and throughput demands of Large Language Models (LLMs) in long contexts. Existing methods typically employ a uniform sparse attention mask, applying the same sparse pattern across different attention heads and input lengths. However, this uniform approach fails to capture the diverse attention patterns inherent in LLMs, ignoring their distinct accuracy-latency trade-offs. To address this challenge, we propose the Mixture of Attention (MoA), which automatically tailors distinct sparse attention configurations to different heads and layers. MoA constructs and navigates a search space of various attention patterns and their scaling rules relative to input sequence lengths. It profiles the model, evaluates potential configurations, and pinpoints the optimal sparse attention compression plan. MoA adapts to varying input sizes, revealing that some attention heads expand their focus to accommodate longer sequences, while other heads consistently concentrate on fixed-length local contexts. Experiments show that MoA increases the effective context length by 3.9×3.9$\times$3.9 × with the same average attention span, boosting retrieval accuracy by 1.57.1×1.5-7.1$\times$1.5 - 7.1 × over the uniform-attention baseline across Vicuna-7B, Vicuna-13B, and Llama3-8B models. Moreover, MoA narrows the capability gaps between sparse and dense models, reducing the maximum relative performance drop from $9\%-36\%$ to within $5\%$ across two long-context understanding benchmarks. MoA achieves a 1.21.4×1.2-1.4$\times$1.2 - 1.4 × GPU memory reduction and boosts decode throughput by $\times5$ 6.7 × for 7B and 13B dense models on a single GPU, with minimal impact on performance.
footnotetext: Corresponding authors: Yu Wang (yu-wang@tsinghua.edu.cn), Xuefei Ning (foxdoraame@gmail.com)
@@ 1 Introduction
Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval dataset. This "Needle In a Haystack" experiment [19] takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions. (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
@@
Large Language Models (LLMs) exhibit remarkable versatility across numerous applications [6, 57, 63]. Central to LLM is the attention mechanism [62], which computes interactions among tokens within a certain span, thereby enabling context understanding. Scaling input length is crucial for enhancing LLM capabilities [7, 60], including fact retrieval, summarization, few-shot learning, question answering and so on [4, 70]. However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges [54, 69, 26, 33].
@@
Previous work proposes sparse attention methods to address the efficiency challenges of long contexts in generative LLMs. These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only [69, 26]. This approach allows the LLM to take long inputs with a fixed attention span, keeping bounded attention computation and KV caching overhead. Following previous works [7, 60], we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold. In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span [18]. Nonetheless, we reveal that uniform masks, like StreamingLLM [69], hardly extend effective context length beyond the span, as shown in Figure 6. Figure 1(b) further illustrates such limitation: with a 50% attention span mask, StreamingLLM fails to accurately retrieve content from the earlier half of the input and performs even worse at longer input lengths. Figure 2 reveals one possible explanation for the problem: while some attention heads focus on local contexts, others encompass the broad span of the entire input sequence. Consequently, the uniform approach fails to achieve a long effective context length as it limits the attention span of the global-context heads, while excessively allocates compute and memory budget for local-context heads. Additionally, as the input length increases, some attention heads need a faster increase in attention span than others to avoid serious performance degradation, as shown in Table 1. Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads. Besides, existing model compression methods [44, 40, 68, 36, 31, 37] use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
@@
  • Calibration Dataset Construction We emphasize the importance of data engineering in LLM compression. Our findings demonstrate that using datasets with long-range dependencies and referencing the original LLM’s responses are crucial for accurately profiling the influences of compression.
@@
Experiments show that MoA achieves $\times5$ to 6.7×6.7$\times$6.7 × throughput improvements on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only 1% average relative degradation in retrieval accuracy. Additionally, MoA achieves over 90% retrieval accuracy with just 25% average density, far surpassing sparse attention baselines that need a density of 75% to 100% for similar performance. On long-context understanding benchmarks, MoA performs comparably to dense models, with a maximum relative performance drop of less than 5%, which is about one-sixth of that observed with the uniform sparse attention baseline. Our code is available at https://github.com/thu-nics/MoA.
@@ 2.1 Attention mechanism
The Multi-Head Self Attention (MHA) mechanism [62] is crucial to the functionality of LLMs. It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections. These matrices, combined with the cached K and V (KV-Cache) from previous sequences, compute the attention matrix (A). This calculation is modified by a causal mask (M) to ensure autoregressive properties, resulting in the output (O), as depicted in Equation 1:
@@ 2.2 Efficient Attention
Efficient methods are proposed to mitigate the computation and memory costs associated with attention. One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage [46, 52, 53, 64, 43, 32] or drop KV-Cache during decode stage [3, 75, 20, 54, 41] based on the input sequences. However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup [52, 64, 43, 25, 24]. Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining [3], or additional accumulated attention score computation [54, 75, 41, 20].
Another branch of work uses static sparse attention, where predefined masks are applied consistently across all processed sentences. Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly. For language understanding models such as BERT [16], various masks are used [73, 5, 9, 76, 69, 26]. But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens [69, 26]. With the local attention pattern, the KV-Cache beyond the current attention span can be dropped, saving much memory for long sequence scenarios. However, the uniform static masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long sequence scenarios. Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling [21, 49, 56, 51, 39, 30, 50, 10, 65]. However, these new mechanisms often require different weights compared to vanilla transformers, imposing significant re-training overhead for LLMs.
Previous works also propose LLM acceleration frameworks [22, 2, 54, 33], as well as kernel-level optimizations [13, 12]. These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
@@ 3 Mixture of Attention (MoA)
Table 1: Retrieval accuracy of Vicuna-7B model using sliding-window sparse attention masks across various model layers, window spans, and input lengths.
@@ 3.1 Mixture of Attention Patterns and Elastic Rules
Heterogeneous Attention Patterns. Different attention heads in LLMs exhibit heterogeneous attention patterns, as shown in Figure 2. For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention. Table 1 demonstrates that applying the same sliding-window sparse attention mask across model layers can lead to a 65% variance in retrieval accuracies. It conforms to the multi-head self-attention design principle of capturing varied information [62], as well as the findings from concurrent research that identifies specific attention heads for global text retrieval [67].
@@ 3.2 Heterogeneous Elastic Rule Search Space
In designing the search space for the MoA mask, we consider the inherently heterogeneous and elastic nature of LLM attention patterns. As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask [5]. Following previous work [69, 26], the initial few tokens (64 tokens for MoA) are not masked. The attention span equals the sliding-window-span plus the number of initial unmasked tokens. For simplicity, we define the attention span S of head $h$ at input length N using a straightforward linear function:
$S_{h}=\alpha_{h}+\beta_{h}\times N,$ (2)
where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length. To ensure hardware-friendliness, we limit the number of distinct rules to at most two per model layer for MoA unless specified otherwise.
@@ 4.1 Attention Influence Profiling
In practice, we use backpropagation on a calibration dataset to calculate the average attention influence $\bar{\textbf{E}}_{h}$ of each head across data items. The average attention influence is calculated respectively for different input lengths. The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch [48]. The detailed calibration dataset setup is discussed in Section 5.
With the average attention influence of each head, MoA can calculate the accuracy loss of applying a candidate elastic rule at a specific input length. The loss is calculated with the sum of masked average attention influence according to the rule. We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 0 and others as 1. We formalize accuracy loss $\Delta L$ as follows:
$\Delta L=\sum_{h}\Delta L_{h,r_{h}}=\sum_{h}\sum_{i}\sum_{j}M_{r_{h},i,j}\bar{E}_{h,i,j}.$ (4)
@@ 4.2 Automatic Optimization
Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi [23]. The detailed problem formulation and solving strategies are discussed in Appendix D.2.
@@ 5 Dataset and Supervision
In this section, we highlight the overlooked importance of calibration dataset design and its supervision objective in LLM compression. Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning [44, 34, 42] and quantization [40, 68, 36, 31]. In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
Current Approach. General language modeling datasets, such as human-written text corpus RedPajama [11], are commonly used as the calibration dataset. These datasets, supervised by next-token-prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens. However, they lack long context dependencies, failing to address the global attention crucial for tasks like long-range retrieval.
Moreover, a notable misalignment exists between the model response and the human-written supervision. Consequently, it leads to inaccuracies when using human responses to compute attention values and gradients during profiling. For example, given the same question, a human might answer ’Blue’, while the model could generate ’The blue color’. Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ’Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ’The’. These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
MoA’s Approach. MoA enhances the calibration dataset by integrating long-range dependencies and model alignment. It utilizes the long-contextual MultiNews dataset [17], which includes summaries heavily dependent on long-range content. The summaries are generated by the original dense model and serve as supervision. Compared to current approaches that adopt human responses as the reference to calculate the loss, using the responses generated by the original model as the supervision can facilitate accurate influence profiling, thus benefiting the compression results.
@@
We show the importance of long-range dependencies by comparing the MoA compression plan generated with different datasets and supervisory methods. In Table 2, RedPajama [11] represents the general language modeling dataset, while MultiNews [17] highlights long-range contexts by aggregating multiple documents on a single incident. Additionally, each MultiNews item includes a human-written summary, providing even stronger long-range dependencies and better performance. Calculating loss on the summary of MultiNews leads to significantly better performance, with a 60% increase in retrieval accuracy and a 0.98 decrease in perplexity.
@@ 6.1 Setups
Baselines. We compare MoA with state-of-the-art static and dynamic sparse attention baselines for LLMs: StreamingLLM [69] and H2O [75]. We define the density of an LLM as the average KV-Cache length divided by the sequence length during the sparse decode stage. Notably, in MoA and StreamingLLM, the KV-Cache length matches the attention span during their sparse prefill stages. In contrast, H2O employs a dense prefill and requires additional computations for dynamic mask generation based on accumulative attention scores.
Models and Benchmarks. We evaluate on vicuna-7b-v1.5-16k, vicuna-13b-v1.5-16k [8] from LMSys, and Llama-3-8B-Instruct-262k [1] from Gradient AI. For effective context length evaluation, we use LongEval [35] to test key-value retrieval accuracy with 100 data items per length level. For comprehensive ability evaluation, we use LV-Eval [70] and LongBench [4], which include 11 and 13 sub-tasks, respectively. For coherence testing, we measure perplexity on four subsets of LongBench [15, 17, 38, 28, 45], representing different capabilities.
MoA Settings. We profile MoA on MultiNews [17] with model summaries at 2k, 4k, and 8k lengths. The optimal compression plan is selected with the validation dataset at 12k. Unless specified, the same plan is used across all benchmarks and lengths for each model. The models are not finetuned.

### v2

@@ Abstract
[…]
footnotetext: Corresponding author: Xuefei Ning (foxdoraame@gmail.com)
@@ 1 Introduction
Figure 1: Retrieval accuracy of the Vicuna-7B model using different attention methods across varying input lengths and retrieval positions on the LongEval benchmark (Li et al., 2023). This retrieval benchmark takes massive key-value pairs as inputs and tests the accuracy to retrieve values based on given keys from diverse positions. (a) Original model with a full attention span; (b) StreamingLLM with half the attention span, showing reduced effectiveness beyond the span; (c) MoA with half the attention span, maintaining effectiveness beyond the span.
@@
Large Language Models (LLMs) exhibit remarkable versatility across numerous applications (Brown et al., 2020; Tay et al., 2022; Wan et al., 2023). Central to LLM is the attention mechanism (Vaswani et al., 2017), which computes interactions among tokens within a certain span, thereby enabling context understanding. Scaling input length is crucial for enhancing LLM capabilities (Chen et al., 2023; Tworkowski et al., 2023), including fact retrieval, summarization, few-shot learning, question answering and so on (Bai et al., 2023; Yuan et al., 2024). However, the ever-growing attention computation and Key-Value Cache (KV-Cache) pose significant efficiency challenges (Sheng et al., 2023; Xiao et al., 2024c; Han et al., 2023; Kwon et al., 2023).
@@
Previous work proposes sparse attention methods to address the efficiency challenges of long contexts in generative LLMs. These methods typically employ a uniform, fixed-span sliding window mask across all heads and input lengths, limiting attention to local contexts only (Xiao et al., 2024c; Han et al., 2023). This approach allows the LLM to take long inputs with a fixed attention span, keeping bounded attention computation and KV caching overhead. Following previous works (Chen et al., 2023; Tworkowski et al., 2023), we quantify the effective context length as the maximum input length where content retrieval accuracy exceeds a 90% threshold. In principle, fixed-span local attention can gradually aggregate global information through multiple model layers, yielding a longer effective context length than each attention span (Feng et al., 2022; Zaheer et al., 2020). Nonetheless, we reveal that uniform masks, like StreamingLLM (Xiao et al., 2024c), hardly extend effective context length beyond the span, as shown in Figure 6. Figure 1(b) further illustrates such limitation: with a 50% attention span mask, StreamingLLM fails to accurately retrieve content from the earlier half of the input and performs even worse at longer input lengths. Figure 2 reveals one possible explanation for the problem: while some attention heads focus on local contexts, others encompass the broad span of the entire input sequence. Consequently, the uniform approach fails to achieve a long effective context length as it limits the attention span of the global-context heads, while excessively allocates compute and memory budget for local-context heads. Additionally, as the input length increases, some attention heads need a faster increase in attention span than others to avoid serious performance degradation, as shown in Table 1. Unfortunately, the uniform approaches do not include heterogeneous rules to scale the attention spans differently for various heads. Besides, existing model compression methods (Men et al., 2024; Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023; Li et al., 2024b) use general language modeling corpora to decide the compression plan, which cannot accurately profile the influence of compression on long-context tasks.
@@
  • Calibration Dataset Construction We emphasize the importance of data engineering in LLM compression. Our findings demonstrate that, instead of relying on general language modeling datasets and human responses, using datasets with long-range dependencies and referencing the original LLM’s responses is essential for accurately profiling the effects of compression.
@@
Experiments show that MoA achieves $\times6$ 8.2 × throughput improvements over FlashAttention2, 1.71.9×1.7-1.9$\times$1.7 - 1.9 × over vLLM framework on 7B and 13B dense LLMs at a 50% density (the average of KV-Cache length / input length), with only a 1% average relative degradation in retrieval accuracy. In Section 6.4, we show that this significant throughput improvements of MoA over FlashAttention2 can be attributed to four factors: (1) the static size of the KV-cache, (2) reduced attention computations, (3) increased batch size enabled by reduced memory usage, and (4) a specialized kernel implementation. Additionally, MoA achieves over 90% retrieval accuracy with just 25% average density, far surpassing sparse attention baselines that need a density of 75% to 100% for similar performance. On long-context understanding benchmarks, MoA performs comparably to dense models, with a maximum relative performance drop of less than 5%, which is about one-sixth of that observed with the uniform sparse attention baseline. Our code is available at https://github.com/thu-nics/MoA.
@@ 2.1 Attention mechanism
The Multi-Head Self Attention (MHA) mechanism (Vaswani et al., 2017) is crucial to the functionality of LLMs. It starts with an input sequence transformed into query (Q), key (K), and value (V) matrices through linear projections. These matrices, combined with the cached K and V (KV-Cache) from previous sequences, compute the attention matrix (A). This calculation is modified by a causal mask (M) to ensure autoregressive properties, resulting in the output (O), as depicted in Equation 1:
@@ 2.2 Efficient Attention
Efficient methods are proposed to mitigate the computation and memory costs associated with attention. One branch of work uses dynamic sparse attention masks to adaptively skip attention computations during prefill stage (Pagliardini et al., 2023; Qu et al., 2022; Roy et al., 2021; Wang et al., 2021; Lu et al., 2021; Kitaev et al., 2020) or drop KV-Cache during decode stage (Anagnostidis et al., 2023; Zhang et al., 2023; Ge et al., 2023; Sheng et al., 2023; Liu et al., 2023a) based on the input sequences. However, due to the complex control and computation flow, dynamic prefill often requires specific hardware to achieve substantial wall-time speedup (Qu et al., 2022; Wang et al., 2021; Lu et al., 2021; Ham et al., 2021; 2020). Additionally, dynamic KV-Cache pruning in the decode stage may require extensive retraining (Anagnostidis et al., 2023), additional accumulated attention score computation (Sheng et al., 2023; Zhang et al., 2023; Liu et al., 2023a; Ge et al., 2023), or extensive memory swap for KV-Cache retrieval (Tang et al., 2024; Xiao et al., 2024a).
Another branch of work uses static sparse attention, where predefined masks are applied consistently across all processed sentences. Thanks to the fixed computation flow, static sparse attention is generally more efficient and GPU-friendly. For language understanding models such as BERT (Devlin et al., 2018), various masks are used (Zaheer et al., 2020; Beltagy et al., 2020; Child et al., 2019; Zhou et al., 2024; Xiao et al., 2024c; Han et al., 2023). But for generative LLMs, the predominant method is the fixed-span sliding window mask with global attention on a few initial tokens (Xiao et al., 2024c; Han et al., 2023). With the local attention pattern, the KV-Cache beyond the current attention span can be dropped, saving much memory for long sequence scenarios. However, the uniform static masks across different attention heads and input lengths are model- and data-agnostic, which can compromise LLMs’ effective context length and lead to suboptimal performance in long sequence scenarios. Our method falls within this category, benefiting from the efficiency and training-free advantages, while addressing the performance limitations encountered by previous methods.
In addition to sparse attention, alternative mechanisms have been proposed to replace traditional attention for long-sequence modeling (Gu & Dao, 2023; Peng et al., 2023; Sun et al., 2023; Poli et al., 2023; Li et al., 2022; Kacham et al., 2023; Peng et al., 2021; Choromanski et al., 2020; Wang et al., 2020). However, these new mechanisms often require different weights compared to vanilla transformers, imposing significant re-training overhead for LLMs.
Previous works also propose LLM acceleration frameworks (Gugger et al., 2022; Aminabadi et al., 2022; Sheng et al., 2023; Kwon et al., 2023), as well as kernel-level optimizations (Dao et al., 2022; Dao, 2023; Shah et al., 2024). These kernel and system optimizations are orthogonal to our work and can be integrated to further enhance efficiency.
@@ 3 Mixture of Attention (MoA)
Table 1: Retrieval accuracy of Vicuna-7B with sliding-window sparse attention across various model layers, window spans, and input lengths.
@@ 3.1 Mixture of Attention Patterns and Elastic Rules
Heterogeneous Attention Patterns. Different attention heads in LLMs exhibit heterogeneous attention patterns, as shown in Figure 2. For example, the first head primarily focuses on local contexts with a narrow-span sliding window, while the third head covers nearly the entire input, indicating global attention. The attention spans of different heads mostly remain constant across various tasks and datasets, as shown in Appendix D.1. Table 1 demonstrates that applying the same sliding-window sparse attention mask across model layers can lead to a 65% variance in retrieval accuracies. It conforms to the multi-head self-attention design principle of capturing varied information (Vaswani et al., 2017), as well as the findings from concurrent research that identifies specific attention heads for global text retrieval (Wu et al., 2024).
@@ 3.2 Heterogeneous Elastic Rule Search Space
In designing the search space for the MoA mask, we consider the inherently heterogeneous and elastic nature of LLM attention patterns. As shown in Figure 3(a), we adopt a hardware-friendly sliding-window mask as our base sparse attention mask (Beltagy et al., 2020). Following previous work (Xiao et al., 2024c; Han et al., 2023), the initial few tokens (64 tokens for MoA) are not masked. The attention span equals the sliding-window-span plus the number of initial unmasked tokens. We define the attention span S of head $h$ at input length N using a straightforward linear function:
$S_{h}=\alpha_{h}+\beta_{h}\cdot N,$ (2)
where $\alpha_{h}$ and $\beta_{h}$ are hyperparameters that control the base span and its expansion rate with input length of a specific attention head.
@@ 4.1 Attention Influence Profiling
In practice, we use backpropagation on a calibration dataset to calculate the average attention influence $\bar{\textbf{E}}_{h}$ of each head across data items. The average attention influence is calculated respectively for different input lengths. The gradient ${\partial L}/{\partial\textbf{A}_{h}}$ is computed using chain derivative in deep learning frameworks like PyTorch (Paszke et al., 2019). The detailed calibration dataset setup is discussed in Section 5.
With the average attention influence of each head, MoA can calculate the accuracy loss of applying a candidate elastic rule at a specific input length. The loss is calculated with the sum of masked average attention influence according to the rule. We denote $\textbf{M}_{r_{h}}$ as the binary mask at head $h$ that corresponds to rule $r$, with masked positions marked as 1 and others as 0. We formalize accuracy loss $\Delta L$ as follows:
$\Delta L=\sum_{h}\Delta L_{h,r_{h}}=\sum_{h}\sum_{i}\sum_{j}M_{r_{h},i,j}\cdot\bar{E}_{h,i,j}.$ (4)
@@ 4.2 Automatic Optimization
Such formulation corresponds to the classic multi-objective mixed-integer-planing problem, which can be effectively solved within minutes using existing linear solvers, like Gurobi (Gurobi Optimization, LLC, 2023). The detailed problem formulation and solving strategies are discussed in Appendix D.3.
@@ 5 Dataset and Supervision
In this section, we highlight the overlooked importance of calibration dataset design and its supervision objective in LLM compression. Calibration datasets are essential for sensitivity analysis across various compression techniques, including weight pruning (Men et al., 2024; Lee et al., 2024; Liu et al., 2023b) and quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024a; Kim et al., 2023). In this work, MoA profiles the attention influence on the calibration dataset, which is crucial for subsequent automatic optimization.
Current Approach. General language modeling datasets, such as human-written text corpus RedPajama (Computer, 2023), are commonly used as the calibration dataset. These datasets, supervised by next-token-prediction on the entire corpus, primarily capture attention patterns coherent with immediately preceding tokens. However, they lack long context dependencies, failing to address the global attention crucial for tasks like long-range retrieval.
Moreover, a notable misalignment exists between the model response and the human-written supervision. Consequently, it leads to inaccuracies when using human responses to compute attention values and gradients during profiling. For example, given the same question, a human might answer ‘Blue’, while the model could generate ‘The blue color’. Using the human answer for supervision, attention influence is inaccurately quantified based on probability shift for predicting ‘Blue’; this diverges from the objective of maintaining crucial attention for the original model prediction, ‘The’. These inconsistencies arise from various factors, including mismatched positions, tones, and synonyms.
MoA’s Approach. MoA enhances the calibration dataset by integrating long-range dependencies and model alignment. Specifically, we utilize the long-contextual MultiNews dataset (Fabbri et al., 2019), which includes summaries heavily dependent on long-range content. The summaries are generated by the original dense model and serve as supervision. Compared to current approaches that adopt human responses as the reference to calculate the loss, using the responses generated by the original model as the supervision can facilitate accurate influence profiling, thus benefiting the compression results.
@@
We show the importance of long-range dependencies by comparing the MoA compression plan generated with different datasets and supervisory methods. In Table 2, RedPajama (Computer, 2023) represents the general language modeling dataset, while MultiNews (Fabbri et al., 2019) highlights long-range contexts by aggregating multiple documents on a single incident. Additionally, each MultiNews item includes a human-written summary, providing even stronger long-range dependencies and better performance. Calculating loss on the summary of MultiNews leads to significantly better performance, with a 60% increase in retrieval accuracy and a 0.98 decrease in perplexity.
@@ 6.1 Setups
Baselines. We compare MoA with state-of-the-art static and dynamic sparse attention methods, including StreamingLLM (Xiao et al., 2024c), InfLLM (Xiao et al., 2024a) and H2O (Zhang et al., 2023). We define the density of an LLM as the ratio of the average in-memory KV-Cache length to the sequence length during the sparse decode stage. Notably, in MoA and StreamingLLM, KV-Cache length equals the attention span during the sparse prefill stage. In contrast, H2O use dense prefill. Besides, H2O and InfLLM require additional computations to dynamically determine the KV-Cache.
Models and Benchmarks. We evaluate on vicuna-{7b, 13b}-v1.5-16k models (Chiang et al., 2023) from LMSys and Llama-3-{8b, 70b}-Instruct-262k models (AI, 2024) from Gradient AI. For long-context retrieval, we use LongEval (Li et al., 2023) to test key-value retrieval accuracy with 100 data items per length level. For long-context understanding, we use LV-Eval (Yuan et al., 2024) and LongBench (Bai et al., 2023), which include 11 and 13 sub-datasets, respectively. For coherence testing, we measure perplexity on four long datasets (Dasigi et al., 2021; Fabbri et al., 2019; Li & Roth, 2002; Hovy et al., 2001; Mohler et al., 2016) with diverse tasks. Unless otherwise specified, performance experiments are restricted to eight A100-80GB GPUs over a 24-hour period, with OOM (Out-Of-Memory) and OOT (Out-Of-Time) conditions noted. Efficiency experiments measure the decode throughput on a single A100-80GB GPU at maximum batch sizes of respective methods.
MoA Settings. We restrict the number of distinct rules to at most two per model layer to ensure inference-time efficiency. We profile MoA on MultiNews (Fabbri et al., 2019) with model summaries at 2k, 4k, and 8k lengths. The optimal compression plan is selected with the validation dataset at 12k. Each model uses the same plan across all benchmarks and lengths. The models are not fine-tuned.
6.2 Accuracy-Throughput Trade-off
Figure 4 shows that MoA advances the Pareto Front in context retrieval accuracy and decode throughput across varied densities. At the same densities, MoA notably enhances throughput by 1.6×1.6$\times$1.6 × to 18.1×18.1$\times$18.1 × compared to H2O and InfLLM, due to its efficient static attention design. The throughput even outperforms StreamingLLM thanks to our customized GPU kernel. Additionally, MoA achieves notably higher retrieval accuracies across all densities. We conduct extensive evaluations for MoA ’s performance and efficiency on various benchmarks across context lengths from 4k to 256k and model sizes ranging from 7B to 70B in subsequent sections.